OpenStudy (anonymous):

The area of a regular hexagon is 30 in.2. Find the length of a side. Round your answer to the nearest tenth. A) 5.9 in. B) 3.4 in. C) 4.5 in. D) 11.5 in. Thanks!

OpenStudy (amistre64):

a hexagon has the property that its inner parts all have the same lengths; its made of equilateral triangles .. 6 of them

OpenStudy (amistre64):

30/6 = and area of 5 for each triangle inside of it

OpenStudy (amistre64):

5 = a^2 sin(60)/2 10 = a^2 (sqrt(3)/2) 10/sqrt(3) = a^2 a = sqrt(10/sqrt(3)) =abt.....

OpenStudy (amistre64):

3.4 after I adjust the 10 to 20 to account for the typing error